OpenStudy (anonymous):

50/3% of a certain sum of money is Rs.250.Find the total sum.

OpenStudy (ganpat):

is that 50 divide by 3% ??

OpenStudy (anonymous):

yeah

OpenStudy (ganpat):

so, the equation says, 50 / 3% of x = 250...

OpenStudy (ganpat):

3% 0f x = 50/250 SO, x * 3/100 = 50 / 250

OpenStudy (anonymous):

culd u jst tell me the annswer

OpenStudy (ganpat):

ha ha, i tried doing that.. and u know what, i got warning for the same :(... So m afraid, cant do that...

OpenStudy (anonymous):

u got warning for wat?

OpenStudy (ganpat):

just telling the answers straight away...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i jst need help as in i am making an answer key to this chapter, so i was in hurry, so i asked for answer!

OpenStudy (anonymous):

nevermind! thnx for the help bye..

OpenStudy (ganpat):

so u not gonna try ??

OpenStudy (anonymous):

plz help!!

OpenStudy (ganpat):

so any tries ?? since yesterday ?? :D

OpenStudy (anonymous):

nope!

OpenStudy (ganpat):

ohkey.. so lets start fresh, what say ?

OpenStudy (anonymous):

hmm.

OpenStudy (anonymous):

plz just calculate and tell me the answer.

OpenStudy (ganpat):

i guess it is not 50/3%... it is (50/3) % ...

OpenStudy (ganpat):

i am afraid dear.. we can do this together..

OpenStudy (anonymous):

OMG I AM REALLY OCCUPIED WITH OTHER THINGS , PLEASE HELP ANYONE IN SOLVING.

mathslover (mathslover):

M i allowed to enter this discussion?

OpenStudy (ganpat):

let the original be x.. SO (50/3)% of x = 250..

OpenStudy (anonymous):

I THINK ANSWER WOULD BE 1500

mathslover (mathslover):

@kajolc.2496 Sorry to say this site only prefers interaction between the asker and the answerer Please Prefer for interaction instead of direct answer: http://openstudy.com/code-of-conduct by the way :YES it is 1500 = x

OpenStudy (anonymous):

IS IT WRIGHT?

OpenStudy (anonymous):

OK I M REALLY SORRY FOR THAT!

mathslover (mathslover):

yes .. according to the solution of the eq. given by @Ganpat

OpenStudy (anonymous):

THANK YOU BOTH!

OpenStudy (ganpat):

@kajolc.2496 : welcome :)